# Visual Evolutionary Optimization on Combinatorial Problems with Multimodal Large Language Models: A Case Study of Influence Maximization

### v1

@@ III-E Complexity Analysis
The complexity of the graph sparsification method centers on several key computations. The betweenness centrality is calculated with a complexity of $O(|V||E|)$. Constructing the subgraph and adjusting edges are achieved in $O(|V^{\prime}|+|E^{\prime}|)$. The process of community merging iterates $(|\mathcal{C}|-|\mathcal{C}^{\prime}|)$ times, where $|\mathcal{C}|$ represents the initial number of communities and $|\mathcal{C}^{\prime}|$ the target count. Each iteration involves identifying the smallest community and calculating its edge connections, requiring $O(|E|)$ time. The running time of MLLM-based evolutionary operators depends on the model’s architecture, with lighter and more efficiently optimized models achieving quicker results.
@@ IV Experimental studies
In the experiment, we examine the effectiveness of MLLM in the evolutionary stages, as well as the sensitivity of the parameters.
@@ IV-A Experimental Settings and Dataset
In the validation, we evaluate our VEO on eight real-world networks obtained from the Network Repository (https://networkrepository.com/). The number of nodes and edges in the simplified networks is fixed at 50 and 100, respectively, to remain manageable for MLLMs. The probability of crossover and mutation are set to 0.2 and 0.1. The size of the initialized population is set to 15 and each strategy is prompted to generate 5 individuals. The reported result is averaged from 20 independent simulations. Communities with fewer than 0.02 of the total nodes will be merged into their nearest, larger community. The initial community distribution of networks is obtained by the FastGreedy algorithm [16]. The number of selected seeds is set to 5 for Netscience and USAir and is set to 10 for the rest of the larger-scale network. The backbone MLLM is gpt-4o-2024-11-20. The visualization settings for different evolutionary operators are listed in Table III.
TABLE II: Basic network information including the number of nodes $|V|$, the number of edges $|E|$, the average degree $\langle K\rangle$, the clustering coefficient $CC$, and the average shortest path length $\langle d\rangle$. $|{\mathcal{C}}|$ and $|{\mathcal{C^{\prime}}}|$ denote the number of communities in the original and simplified networks.
  Network $|V|$ $|E|$ $\langle K\rangle$ $CC$ $\langle d\rangle$ $|{\mathcal{C}}|$ $|{\mathcal{C^{\prime}}}|$
@@ IV-C Examination of MLLM as Initialization Operator
USAir 48.84$\pm$2.23(-) 49.27$\pm$1.57(-) 49.12$\pm$1.88($\approx$) 49.19$\pm$1.51($\approx$) 49.92$\pm$1.98
Netscience 16.24$\pm$0.56(-) 15.89$\pm$0.80(-) 16.40$\pm$0.53($\approx$) 16.35$\pm$0.41(-) 16.63$\pm$0.26
Polblogs 183.32$\pm$13.11(-) 213.77$\pm$5.94($\approx$) 203.93$\pm$8.09(-) 206.54$\pm$8.04(-) 216.57$\pm$7.43
Facebook 236.46$\pm$49.69(-) 407.45$\pm$8.86($\approx$) 380.77$\pm$31.13(-) 376.12$\pm$36.48(-) 410.57$\pm$5.80
WikiVote 315.58$\pm$28.19(-) 545.13$\pm$15.67(-) 498.45$\pm$25.11(-) 485.01$\pm$22.03(-) 563.67$\pm$10.94
Rutgers89 327.80$\pm$72.36(-) 778.96$\pm$25.65(-) 692.16$\pm$44.54(-) 671.41$\pm$34.45(-) 790.93$\pm$8.36
MSU24 363.87$\pm$102.34(-) 1152.17$\pm$27.81(-) 1098.03$\pm$52.09(-) 1033.17$\pm$113.48(-) 1170.79$\pm$15.29
Texas84 543.29$\pm$147.81(-) 2594.51$\pm$144.18($\approx$) 2063.10$\pm$208.67(-) 2109.58$\pm$190.13(-) 2629.62$\pm$80.04
[…]
Avg ranking 4.88 2.38 3.25 3.50 1.00
To compare different methods, we use the Wilcoxon test with a 95% confidence interval to statistically examine their difference. The ‘+’ indicates statistically better performance, $\approx$’ indicates no significant difference, and ‘-’ indicates statistically worse performance. Table IV-C presents the optimization results regarding fitness values achieved through different initialization modes across eight networks. The fitness values are reported as averages with standard deviations. The MLLMs method consistently outperforms other modes across all networks, as indicated by the highest fitness values and its average ranking of 1.00, demonstrating its superiority. This analysis highlights the superior capability of MLLMs in initializing the population for evolutionary optimization more effectively compared to traditional approaches. In addition, the results show that Refined Random initialization (sampled from the simplified network) outperforms the other three non-MLLM initialization methods (derived from the original network) in 7 out of 8 cases except for Netscience, demonstrating the effectiveness of our graph sparsification approach.
  Graph F-statistic p-value Different
@@
In addition, we investigate the accuracy of MLLMs in following instructions. Specifically, we examine whether the initial population obtained through various strategies exhibits any differences. Table V presents the ANOVA results for three different agent-based initialization methods used in various networks. These agents operate under strategies formulated by MLLMs, which demonstrate an acute awareness of network structural topologies and can execute tasks according to specific instructions. The results indicate significant differences in the performance of these agents in seven out of the eight networks examined, with only MSU24 and Texas84 showing no significant differences among the agents’ strategies. This result also demonstrates the spatial intelligence of MLLMs and their potential in the graph problems.
@@
To gain a perspective on the node selection behavior of MLLMs during mutation, we compared the degree distribution of nodes suggested for removal and addition by MLLMs in Figure 2. As seen, the average degree of nodes suggested for addition is clearly higher than those suggested for removal, indicating a strategy favoring the addition of more highly connected nodes to maximize influence or connectivity potentially. The marked contrast in the degree distributions, especially in networks, indicates the MLLMs’ awareness of the network structure and the ability to intelligently identify critical nodes, thereby enhancing the overall strategy for influence maximization.
@@
USAir 44.03$\pm$3.17($\approx$) 43.46$\pm$2.62(-) 43.74$\pm$2.91(-) 45.02$\pm$2.49
Netscience 14.04$\pm$1.11(-) 14.55$\pm$1.15 ($\approx$) 14.75$\pm$0.94($\approx$) 14.63$\pm$1.13
Polblogs 187.95$\pm$8.90($\approx$) 193.65$\pm$10.50($\approx$) 191.44$\pm$8.82($\approx$) 190.75$\pm$6.94
Facebook 354.58$\pm$36.72(-) 353.31$\pm$42.28($\approx$) 356.00$\pm$34.24($\approx$) 367.90$\pm$27.03
WikiVote 480.32$\pm$20.59(-) 487.36$\pm$24.42(-) 501.01$\pm$22.15($\approx$) 506.40$\pm$22.91
Rutgers89 686.99$\pm$39.65(-) 698.56$\pm$46.61($\approx$) 718.94$\pm$44.73($\approx$) 710.39$\pm$45.16
MSU24 1017.86$\pm$140.14(-) 1080.24$\pm$339.29($\approx$) 1071.14$\pm$39.23($\approx$) 1084.14$\pm$39.20
Texas84 2197.38$\pm$225.31(-) 2339.35$\pm$168.78($\approx$) 2269.19$\pm$172.68($\approx$) 2277.91$\pm$213.75
@@ IV-G Computational Cost Analysis
gpt-4o-2024-11-20 2.46$\pm$0.52 1.75$\pm$0.41
Gemeni-2.0-flash-lite 2.28$\pm$0.46 2.18$\pm$0.45
Claude-3.7-sonnet 5.59$\pm$1.34 4.11$\pm$1.30
Qwen-vl-max 3.11$\pm$0.95 1.81$\pm$0.56
@@
In our work, we focus on MLLM-based optimization of combinatorial problems, where a critical aspect is ensuring that the model effectively perceives and reasons over the underlying network structure. To achieve this, the majority of token usage is dedicated to depicting the network, either visually (via images) or textually to provide the LLMs or MLLMs with sufficient structural context. In contrast, the actual prompt instructions used for optimization are minimal and static across tasks, contributing only marginally to the overall token cost. Therefore, our cost analysis specifically isolates and compares the cost of network depiction in image and text formats, which represents the dominant share of token usage. Figure 8 illustrates how the token cost scales with network size under different input representations for both LLM- and MLLM-based models. For LLM inputs (Adjacency and Incident methods) [52], token counts grow significantly as the number of edges or nodes increases, highlighting a direct linear relationship between network size and input complexity. In contrast, for MLLM inputs based on images, the token cost remains almost constant regardless of the network scale, with differences between the 400×400 and 768×768 pixel image sizes. This indicates that visual representations via images offer much greater scalability for handling large networks, as they decouple the input size from the underlying network complexity.
@@ IV-H Generalizability Test
To investigate the generalizability of VEO, we apply it to the network dismantling problem on the Karate network [64], as shown in Figure 9. In this task, fitness is measured as the difference between the original network size and the size of the largest connected component after node removals. Note that network dismantling is fundamentally different from influence maximization, as the latter focuses on maximizing widespread reach, whereas dismantling aims to disrupt connectivity by targeting structurally critical nodes. The results show that the MLLM-based method consistently outperforms the standard evolutionary strategy across iterations, achieving higher fitness values and faster convergence. The descriptions on the right illustrate how MLLMs are prompted to perform task-specific crossover and mutation operations, confirming the adaptability of the framework to different optimization objectives.
@@ V-A Representation-Level Enhancements
  • Graph Reduction: Techniques such as sparsification or abstraction could be explored to reduce graph complexity while preserving essential relational patterns, enabling MLLMs to operate more efficiently on large-scale graphs.

### v2

@@ III-E Complexity Analysis
[…]
@@ IV Experimental studies
In the experiment, we examine the effectiveness of MLLM in the evolutionary stages, as well as the parameter sensitivity.
@@ IV-A Experimental Settings and Dataset
In the validation, we evaluate our VEO on eight real-world networks obtained from the Network Repository (https://networkrepository.com/). The number of nodes and edges in the simplified networks is fixed at 50 and 100, respectively, to remain manageable for MLLMs. The probability of crossover and mutation are set to 0.2 and 0.1. The size of the initialized population is set to 15 and each strategy is prompted to generate 5 individuals. The reported result is averaged from 20 independent simulations. Communities with fewer than 2% of the total nodes will be merged into their nearest, larger community. The initial community distribution of networks is obtained by the FastGreedy algorithm [16]. The number of selected seeds is set to 5 for Netscience and USAir and is set to 10 for the rest of the larger-scale network. The backbone MLLM is Gpt-4o-2024-11-20. The visualization settings for different evolutionary operators are listed in Table III.
TABLE II: Basic network information including the number of nodes $|\text{V}|$, the number of edges $|\text{E}|$, the average degree $\langle\text{K}\rangle$, the clustering coefficient $C$, and the average shortest path length $\langle\text{d}\rangle$. $|{\mathcal{C}}|$ and $|{\mathcal{C^{\prime}}}|$ denote the number of communities in the original and simplified networks.
  Network $\mathbf{|V|}$ $\mathbf{|E|}$ $\mathbf{\langle K\rangle}$ $\mathbf{CC}$ $\mathbf{\langle d\rangle}$ $\mathbf{|{\mathcal{C}}|}$ $\mathbf{|{\mathcal{C^{\prime}}}|}$
@@ IV-C Examination of MLLM as Initialization Operator
[…]
Avg ranking 4.88 2.38 3.25 3.50 1.00
To compare different methods, we use the Wilcoxon test with a 95% confidence interval to statistically examine their difference. The ‘+’ indicates statistically better performance, $\approx$’ indicates no significant difference, and ‘-’ indicates statistically worse performance. Table IV-C presents the optimization results regarding fitness values achieved through different initialization modes across eight networks. The fitness values are reported as averages with standard deviations. The MLLM-based method consistently outperforms other modes across all networks, as indicated by the highest fitness values and its average ranking of 1.00, demonstrating its superiority. This analysis highlights the superior capability of MLLMs in initializing the population for evolutionary optimization more effectively compared to traditional approaches. In addition, the results show that Refined Random initialization (sampled from the simplified network) outperforms the other three non-MLLM initialization methods (derived from the original network) in 7 out of 8 cases except for Netscience, demonstrating the effectiveness of our graph sparsification approach.
  Graph F-statistic P-value Different
Netscience 26.35 $2.92\text{E-}8$ Yes
USAir 8.42 $2.76\text{E-}4$ Yes
Polblogs 4.90 $8.00\text{E-}3$ Yes
Facebook 3.56 $2.96\text{E-}2$ Yes
WikiVote 5.53 $4.36\text{E-}3$ Yes
Rutgers89 3.30 $3.80\text{E-}2$ Yes
MSU24 1.24 $2.91\text{E-}1$ No
Texas84 2.03 $1.33\text{E-}1$ No
In addition, we investigate the accuracy of MLLMs in following instructions. Specifically, we examine whether the initial population obtained through various strategies exhibits any differences. Table V presents the ANOVA results for three different agent-based initialization methods used in various networks. These agents operate under strategies formulated by MLLMs, which demonstrate an acute awareness of network structural topologies and can execute tasks according to specific instructions. The results indicate significant differences in the performance of these agents in 6 out of 8 networks examined, with only MSU24 and Texas84 showing no significant differences among the agents’ strategies. This result also demonstrates the spatial intelligence of MLLMs and their potential in the graph problems.
@@
To gain a perspective on the node selection behavior of MLLMs during mutation, we compared the degree distribution of nodes suggested for removal and addition by MLLMs in Figure 2. As seen, the average degree of nodes suggested for addition is clearly higher than those suggested for removal, indicating a strategy favoring the addition of more highly connected nodes to maximize influence or connectivity potentially. The marked contrast in the degree distributions indicates the MLLMs’ awareness of the network structure and the ability to intelligently identify critical nodes, thereby enhancing the overall strategy for influence maximization.
@@
[…]
@@ IV-G Computational Cost Analysis
[…]
@@
In our work, we focus on MLLM-based optimization of combinatorial problems, where a critical aspect is ensuring that the model effectively perceives and reasons over the underlying network structure. To achieve this, the majority of token usage is dedicated to depicting the network, either visually (via images) or textually to provide the LLMs or MLLMs with sufficient structural context. In contrast, the actual prompt instructions used for optimization are minimal and remain static across tasks, accounting for only a marginal portion of the overall token cost. Therefore, our cost analysis specifically isolates and compares the cost of network depiction in image and text formats, which represents the dominant share of token usage. Figure 8 illustrates how the token cost scales with network size under different input representations for both LLM- and MLLM-based models. For LLM inputs (Adjacency and Incident methods) [52], token counts grow significantly as the number of edges or nodes increases, highlighting a direct linear relationship between network size and input complexity. In contrast, for MLLM inputs based on images, the token cost remains almost constant regardless of the network scale, with differences between the 400×400 and 768×768 pixel image sizes. This indicates that visual representations via images offer much greater scalability for handling large networks, as they decouple the input size from the underlying network complexity.
@@ IV-H Generalizability Test
To investigate the generalizability of VEO, we apply it to the network dismantling problem on the Karate network [64], as shown in Figure 9. In this task, fitness is defined as the difference between the size of the original network and the size of the largest connected component after node removal. Note that network dismantling is fundamentally different from influence maximization, as the latter focuses on maximizing widespread reach, whereas dismantling aims to disrupt connectivity by targeting structurally critical nodes. The results show that the MLLM-based method consistently outperforms the standard evolutionary strategy across iterations, achieving higher fitness values and faster convergence. The descriptions on the right illustrate how MLLMs are prompted to perform task-specific crossover and mutation operations, confirming the adaptability of the framework to different optimization objectives.
@@ V-A Representation-Level Enhancements
  • Graph Reduction: Techniques such as sparsification and abstraction can be further explored to reduce graph complexity while more accurately preserving essential relational patterns, thereby enabling MLLMs to operate more efficiently on large-scale graphs.